# Semantic-ICP: Iterative Closest Point for Non-rigid Multi-Organ Point Cloud Registration

### v2

@@ Abstract
Point cloud registration is important in computer-aided interventions (CAI). While learning-based point cloud registration methods have been developed, their clinical application is hampered by issues of generalizability and explainability. Therefore, classical point cloud registration methods, such as Iterative Closest Point (ICP), are still widely applied in CAI. ICP methods fail to consider that: (1) the points have well-defined semantic meaning, in that each point can be related to a specific anatomical label; (2) the deformation required for registration needs to follow biomechanical energy constraints. In this paper, we present a novel semantic ICP (SemICP) method that handles multiple point labels and uses linear elastic energy regularization. We use semantic labels to improve the robustness of the closest point matching and propose a novel point cloud deformation representation to apply explicit biomechanical energy regularization. Our experiments on a trans-oral robotic surgery ultrasound-computed tomography registration dataset and two public Learn2reg challenge datasets show that our method improves the Hausdorff distance and mean surface distance compared with other point-matching-based registration methods.
{IEEEkeywords}
@@ 1 Introduction
[…]
Registration is an important problem in computer-aided intervention (CAI). To enhance visualization of surgical planning, CAI systems require registration of preoperative volumes to intraoperative volumes, meshes, or surfaces. Point clouds describe critical anatomical structures as spatial points, providing a modality-agnostic, sparse and flexible representation. Point clouds also contain less sensitive information that can identify patients than dense volumes, such as the metadata embedded in DICOM files, addressing privacy concerns in medical practice. Therefore, point cloud registration has been applied in different CAI systems [1].
Deep learning models have emerged in general point cloud registration [2, 3], but they are data-hungry and struggle to generalize to different tasks. A recent study shows that the deep learning model trained on general point clouds cannot be applied to CT patient registration out-of-the-box and has lower performance compared with RANSAC initialization and ICP refinement [4]. Finetuning deep learning models on specific datasets is required to achieve a preferable performance [5]. However, this may not be possible due to the limited training data in a specific CAI application. Another concern regarding deep learning models is that they operate as a black-box, which is suboptimal for medical procedures since explainability can enhance both clinician and patient trust in CAI systems. A generalizable and explainable optimization-based point cloud registration method, therefore, is more preferable in CAI.
One of the most representative optimization-based registration algorithms, Iterative Closest Point (ICP) [6], is widely used in CAI [1] and has been applied in recent navigation systems for spine surgery [7] and oral and maxillofacial surgery [8]. Compared with Gaussian Mixture Models-based registrations [9], point matching-based ICP has a lower computation time, making it preferable in intervention procedures. However, ICP was originally developed for general computer vision that did not utilize medical domain knowledge. First, the semantic information of the point cloud, such as the points’ anatomical labels, is not utilized. Second, non-rigid ICP-based registration [10, 11, 12] regularizes the deformation by local affine regularization, ignoring the underlying biomechanical effects. Addressing these problems, we propose a novel and generalizable point-matching registration method that utilizes the semantic information and introduces explicit linear biomechanical energy regularization.
@@ 1.1.1 ICP and its Variants
ICP [6] was developed over 30 years ago, but it is still widely applied in medical point cloud registration for CAI. ICP iteratively finds the closest point pairs in two point clouds, and the corresponding point pairs are used to estimate a rigid transformation to move the source point cloud closer to the target point cloud. The initial alignment between the two point clouds should be close enough to ensure convergence, making ICP sensitive to initialization and outliers, and it tends to get trapped in local minima. To alleviate problems with initialization, Global-ICP [13] performs a global search over 3D space. Fast-robust-ICP [14] uses Welsch’s function to improve the robustness of the error metric. However, these methods estimate rigid transformations, which cannot correct for point cloud deformation, and therefore their utility is limited in medical applications.
To address non-rigid registration, non-rigid ICP [10] extends ICP to estimate the surface deformation using a similar nearest-point search to find the point correspondences. The deformation is regularized using a local affine transformation. It has been used in ultrasound (US)-guided augmented reality navigation for liver procedures [15]. More recent non-rigid ICP registration methods focus on accelerating convergence and increasing robustness. Yao et al. [11] use a deformation graph with a globally smooth robust estimator to regularize the deformation, and then use accelerated optimization to improve the runtime [12]. However, these methods are not regularized by physical parameters; they only consider local affine regularization, which cannot capture the underlying biomechanical characteristics within medical point clouds.
@@ 1.1.2 Semantic-assisted Point Cloud Registration
Using segmentation for registration is common in 3D volume registration [16, 17, 18]. Medical images contain rich semantic information, including tissue and organ types, which can be leveraged to improve region matching accuracy. In contrast, using semantic segmentation in point cloud registration remains relatively underexplored, likely due to the difficulty of segmenting traditional point clouds. Zaganidis et al. [19] use PointNet to segment light detection and ranging (LiDAR) point clouds and use the segmentation to aid point matching. Wang et al. [20] segment LiDAR point cloud to cars or the background before performing ICP. However, they use rigid registration, which cannot account for surface deformation. Point clouds derived from medical images can be segmented using various techniques, including deep learning [21], or through novel imaging modalities such as T-mode™ ultrasound, which can identify tissue type automatically [22]. Several studies have explored incorporating segmentation into medical point cloud registration. Hiep et al. [23] segment bones and arteries from US and perform ICP registration on them separately. Thomson et al. [16] segment the hepatic and portal veins, and use Coherent Point Drift to register the vessels independently. These studies do not account for interactions between organs or surrounding tissues, failing to capture their interdependent motion. Jiang et al. [24] use segmentation to align US and computed tomography (CT) of the ribs, and the non-rigid point cloud refinement relies on the specific graph structure of cartilage. A generalizable non-rigid registration method that does not rely on specific structures has a broader application.
@@ 1.1.3 Biomechanics-inspired Regularization in Point Cloud Registration
Previous work introducing biomechanical constraints into point cloud registration requires sophisticated finite element modeling (FEM). The regularization can be introduced through iterative FEM directly [25, 26, 27], but FEM is time-consuming and computationally intensive. Data-driven biomechanical models implicitly introduce the constraints by learning from finite element simulation [28, 29, 30], but the trained models lack generalization to new data. Physics or biomechanics-inspired constraints have been introduced through a discrete approximation to the continuous elastic energy in 3D voxels [31, 32, 33]. However, the challenge of introducing this regularization into point cloud registration is that the point clouds only represent the deformation field as surface point movement, which solely describes the boundary conditions of the deformation field. Recently, Heinrich et al. [34, 35] introduced voxelized regularization in point cloud registration, applying spatial B-spline smoothing in voxelized 3D space to stabilize deep learning-based point cloud registration, though this voxelized concept has not been used in optimization-based point cloud registration. Their work also does not consider explicit biomechanical energy regularization, which is necessary for medical point cloud registration.
@@ 1.2 Contributions
In this paper, we propose a new point-matching-based registration method for medical point cloud registration that addresses the lack of semantic information and biomechanical energy regularization. We first propose using semantic labels in the point-matching step, then introduce a novel deformation representation in point clouds and explicit biomechanical energy regularization into point cloud registration. The method was tested on three entirely different datasets and achieved the best surface matching performance compared with other point-matching-based registration algorithms. To the best of our knowledge, our contributions include: (1) the first work to systematically use semantic labels for non-rigid point cloud alignment; (2) the first work to use explicit biomechanical energy regularization in point cloud registration; (3) validation of the effectiveness of (1) and (2) in medical point clouds, with a fixed set of hyperparameters that underwent minimal tuning.
The rest of this paper is organized as follows. Section 2 describes the proposed method, and Section 3 presents validation results on three datasets, ablation studies, and sensitivity analyses. Section 4 provides the discussion.
@@ 2 Methods
We define the source point cloud as $P=\{\bm{p_{0}},\bm{p_{1}},...,\bm{p_{N}}\}$ and the target as $Q=\{\bm{q_{0}},\bm{q_{1}},...,\bm{q_{M}}\}$, with the points $\bm{p},\bm{q}\in\mathbb{R}^{3}$. The point clouds are associated with semantic labels $L_{p}=\{x_{0},x_{1},...,x_{N}\}$ and $L_{q}=\{y_{0},y_{1},...,y_{M}\}$. The labels belong to the label set $L=\{l_{1},l_{2}...,l_{K}\}$, where $K$ is the number of labels. We assume that the observed points have at least two labels, $K\geq 2$. Apart from the 3D location of the points, the point clouds also have local normals at each point: $N_{p}=\{\bm{n_{0}},...,\bm{n_{N}}\}$, $N_{q}=\{\bm{m_{0}},...,\bm{m_{M}}\}$. The local normals can be estimated using surface reconstruction methods such as ball pivoting [36]. The registration workflow of Semantic-ICP (SemICP), is shown in Fig. 1; it first estimates a rigid transform $\bm{T}\in SE(3)$, then refines it by non-rigid registration.
@@ 2.1 Rigid Initialization
The optimal rigid transform $\bm{T}\in SE(3)$ is iteratively estimated by minimizing a cost function that measures the difference between point clouds, as described in Algorithm 1. SemICP uses the semantic information during point pair matching, such that matched point pairs have the same semantic labels. At each iteration $i$, for each point $\bm{p_{j}}$ from $P$, we find the closest point $\bm{q_{j}}$ in $Q$ that has the smallest L2 distance with the same semantic label as $\bm{p_{j}}$. KD-tree is used to accelerate the search for matched point pairs. The update value $\Delta\bm{T}^{i}$ is automatically estimated using the autograd function in PyTorch. We use an Adam optimizer [37] to minimize the cost function $L_{rigid}$ that measures the point-to-plane distance [38] of matched point pairs at each iteration:
@@
6:    Acquire matched $\bm{q_{j}}^{i}$ with smallest L2 distance.
@@ 2.2 Non-rigid Registration
$\bm{d}_{j}$ is the surface point deformation at point $\bm{p}_{j}$. We compute $\bm{d}_{j}$ using a trilinear interpolation of the deformation at the control points that are in $\bm{p}$’s neighborhood.
@@
We use linear elastic regularization [39, 33] to regularize the deformation. In continuum mechanics, the linear elastic energy can be described using the Navier-Lamé Equation. We approximate the Navier-Lamé Equation with the numerical differentiation at the control points and add a scaling factor to normalize the energy, where $\Delta$ is the step size of the control points.
@@
In addition to the above energy regularization, the regularization term also includes magnitude regularization $Reg_{mag}$ to reduce unwanted large deformation, and a gradient regularization $Reg_{grad}$ to encourage control grid smoothness:
@@ 3.1 Datasets
The experiments are performed on one private dataset derived from trans-oral robotic surgery (TORS), and two public datasets from Learn2Reg challenges [40]. We use the training data in the Learn2Reg challenges, since label maps for test data are not included in the public dataset link [^1] https://learn2reg.grand-challenge.org/Datasets/. 3D slicer is used to extract point clouds from the label maps [41].
TORS: This dataset contains 3D freehand US and CT images of the neck collected from patients who underwent TORS for oral cancer between January 2022 and November 2024 at the Vancouver General Hospital (Vancouver, BC, Canada). This study received ethics approval from the University of British Columbia Clinical Research Ethics Board (H19-04025). Informed consent was obtained from all participants included in this study. The procedures used in this study adhere to the tenets of the Declaration of Helsinki. Diagnostic CT data were collected preoperatively, and US data were collected before surgery while patients were anesthetized. A BK3500 with a 14L3 linear 2D transducer (BK Medical, Burlington, MA) and a Polaris Spectra (Northern Digital, ON, Canada) were used to collect freehand 3D US. PLUS [42] was used to calibrate the US probe and collect tracked US data. The image depth was set to 4 or 5 cm (depending on patient anatomy) at 9 MHz, and US imaging was carried out ipsilateral to the tumor. A medical student and an experienced research assistant labeled the carotid artery, jugular vein, and larynx/mandibular bone in 2D US frames and CT volumes, then the 2D US label maps were reconstructed in 3D. The dataset contains 24 pairs of US-CT images from 12 patients. The bifurcation point and the axis of the carotid in US and CT were used to pre-align volumes.
AbdominalMRCT [43]: The Learn2reg dataset contains 8 pairs of intra-patient abdominal magnetic resonance (MR) and CT images. The label maps include 3D segmentation of 4 anatomical labels of liver, spleen, right kidney and left kidney.
HippocampusMR [44]: The Learn2reg challenge dataset includes MR images from 90 healthy adults and 105 adults with a non-affective psychotic disorder, along with segmentations of the head and body of the hippocampus. We use the validation registration pairs provided in the dataset’s JSON files, resulting in 60 inter-patient registration pairs.

### v3

@@ Abstract
Point cloud registration is important in computer-aided interventions (CAI). While learning-based point cloud registration methods have been developed, their clinical application is hampered by issues of generalizability and explainability. Therefore, classical point cloud registration methods, including Iterative Closest Point (ICP), are still widely applied in CAI. ICP methods fail to consider that: (1) the points have well-defined semantic meaning, in that each point can be related to a specific anatomical label; (2) the deformation required for registration needs to follow biomechanical energy constraints. In this paper, we present a novel non-rigid semantic ICP (SemICP) method that handles multiple point labels and uses linear elastic energy regularization. We use semantic labels to improve the robustness of closest point matching and propose a novel point cloud deformation representation that incorporates explicit biomechanical energy regularization. Our experiments on four datasets show that our method significantly improves the Hausdorff distance and mean surface distance compared with other point cloud registration methods. We also demonstrate that integrating deep learning segmentation models with our registration pipeline enables effective alignment of US and MR point clouds.
keywords:
@@
journal: Medical Image Analysist1t1footnotetext: The work is supported by an NSERC Discovery Grant and the Charles Laszlo Chair in Biomedical Engineering held by Dr. Salcudean, the VCHRI Innovation and Translational Research Awards, the University of British Columbia Department of Surgery Seed Grant and CIHR Project Grant held by Dr. Prisman.
[…]
[label1]organization=Department of Electrical and Computer Engineering, University of British Columbia, addressline=2332 Main Mall, city=Vancouver, postcode=V6T 1Z4, state=BC, country=Canada
[…]
[label2]organization=Faculty of Medicine, University of British Columbia, addressline=317 – 2194 Health Sciences Mall, city=Vancouver, postcode=V6T 1Z3, state=BC, country=Canada
@@ 1 Introduction
Registration is an important problem in computer-aided intervention (CAI). To enhance visualization of surgical planning, CAI systems require registration of preoperative volumes to intraoperative volumes, meshes, or surfaces. Point clouds represent critical anatomical structures as spatial points, providing a modality-agnostic, sparse and flexible representation. In particular, in ultrasound (US) to cross-sectional image registration problems, US typically provides only partial observations of internal structures, making volume-based registration challenging. Moreover, bone surfaces and other highly reflective structures in US cause acoustic shadows and incomplete volumes. As a result, converting segmented images to point clouds has become a common and effective approach for US registration in US-guided CAI systems (Lei et al. (2023); Jiang et al. (2025); Hiep et al. (2025)).
Deep learning models have emerged in point cloud registration (Wang and Solomon (2019); Poiesi and Boscaini (2022)), but they are data-hungry and struggle to generalize to different tasks. Indeed, a recent study shows that a deep learning model trained on general point clouds cannot be applied to CT patient registration out-of-the-box and has lower performance compared with RANSAC initialization and ICP refinement (Weber et al. (2024)). In general, fine-tuning deep learning models on specific datasets is required to achieve adequate performance (Zhang et al. (2024)). However, this is rarely feasible due to the limited training data in a specific CAI application. Furthermore, deep learning models often lack the explainability required for clinicians to trust their application in medical interventions. Therefore, a generalizable and explainable optimization-based point cloud registration is preferable in CAI.
One of the most representative optimization-based registration algorithms, Iterative Closest Point (ICP) (Besl and McKay (1992)), is widely used in CAI (Ma et al. (2023)) and has been applied in recent navigation systems for spine surgery (Liebmann et al. (2024)) and oral and maxillofacial surgery (Bayrak et al. (2020)). However, the semantic information of the point cloud, such as the points’ anatomical labels, is not utilized in the original ICP. Furthermore, most registration tasks in medical imaging involve soft tissue, requiring non-rigid ICP-based registration. Prior methods deal with deformation by locally affine regularization, ignoring the underlying biomechanical effects (Amberg et al. (2007); Yao et al. (2020, 2023)). Addressing these problems, we propose and validate a novel non-rigid ICP registration method that utilizes the semantic information and introduces explicit linear biomechanical energy regularization.
@@ 1.1.1 ICP and its Variants
ICP (Besl and McKay (1992)) was developed over 30 years ago, but it is still widely applied in medical point cloud registration for CAI. ICP iteratively finds the closest point pairs in two point clouds, and the corresponding point pairs are used to estimate a rigid transformation to move the source point cloud closer to the target point cloud. The initial alignment between the two point clouds should be close enough to ensure convergence, making the original ICP algorithm sensitive to initialization and outliers and less effective due to local minima. To alleviate these problems, Global-ICP (Yang et al. (2015)) performs a global search over 3D space. Fast-robust-ICP (Zhang et al. (2022)) uses Welsch’s function to improve the robustness of the error metric. However, the above ICP methods estimate only rigid transformations. Therefore, their utility is limited to medical applications in which tissue deformation is negligible.
To address non-rigid registration, non-rigid ICP (Amberg et al. (2007)) extends ICP to estimate the surface deformation using a similar nearest-point search to find the point correspondences. The deformation is regularized using a local affine transformation. It has been used in US-guided augmented reality navigation for liver procedures (Ma et al. (2022)). More recent non-rigid ICP registration methods focus on accelerating convergence and increasing robustness. Yao et al. (2020) use a deformation graph with a globally smooth robust estimator to regularize the deformation, and then use accelerated optimization to improve the runtime (Yao et al. (2023)). However, these methods are not regularized by physical parameters; they only consider local affine regularization, which cannot capture the underlying biomechanical characteristics within medical point clouds.
@@ 1.1.2 Semantic-assisted Point Cloud Registration
To our best knowledge, incorporating semantic segmentation in point matching for ICP has only been presented in the context of rigid registration in autonomous driving, showing improved accuracy compared with regular ICP. Zaganidis et al. (2018) use PointNet to segment light detection and ranging (LiDAR) point clouds and use the segmentation to aid point matching. Wang et al. (2020) segment the LiDAR point cloud to cars or the background before performing ICP. For medical applications, we became aware of concurrent work published after our preprint (Flepp et al. (2025)) that integrates point label-informed matching in bone CT-X-ray rigid registration. However, their work is limited to rigid registration and therefore cannot account for surface deformation, which is critical in medical point cloud registration. Our approach remains distinct in its focus on non-rigid registration and in providing a more systematic evaluation across four datasets, whereas the concurrent work was evaluated on only one dataset. Previous work that utilized the semantic information of point labels performed registration on each anatomy separately. Hiep et al. (2025) segment bones and arteries from US and perform ICP registration on them separately. However, these studies do not account for interactions between organs or surrounding tissues, failing to capture their interdependent motion. Our method introduces a novel and practical control grid-based deformation representation to model inter-object interaction with a biomechanical energy-inspired regularizer in the point cloud registration step. Soft semantic-based matching that learns feature expressions of the point for point matching has also been proposed (Hansen et al. (2019); Mekhzoum et al. (2024)). However, the learned features can be difficult to generalize to different anatomies, while utilizing the consistency of the anatomical labels from the upstream tasks can be applied to any segmentation-generated point clouds, without requiring an anatomy-specific feature extraction step.
@@ 1.1.3 Biomechanics-inspired Regularization in Point Cloud Registration
Previous work introducing biomechanical constraints into point cloud registration requires sophisticated finite element modeling (FEM). The regularization can be introduced through iterative FEM directly (Khallaghi et al. (2015); Ringel et al. (2023); Yang et al. (2025)), but FEM is time-consuming and computationally intensive. Data-driven biomechanical models implicitly introduce the constraints by learning from finite element simulation (Fu et al. (2021a, b); Wang et al. (2024)). However, these models have only been evaluated on the organs used for training, making it unclear whether they can generalize to other organs. Retraining the models for a new domain, on the other hand, requires generating simulated data with time-consuming FEM. Physics or biomechanics-inspired regularization can improve physical plausibility, and it has been introduced through a discrete approximation to the continuous elastic energy in 3D voxels (Broit (1981); Fischer and Modersitzki (2008); Nir et al. (2013)). However, the challenge of introducing this regularization into point cloud registration is that the point clouds only represent the deformation field as surface point movement, which solely describes the boundary conditions of the deformation field. Recently, Heinrich et al. (2023, 2025) introduced novel regularization in point cloud registration by voxelizing the point clouds to a dense representation to stabilize deep learning-based point cloud registration. Compared with dense voxels, we propose to use control point–based deformation that works directly on the points to reduce the memory cost. Their work also does not consider explicit biomechanical energy regularization, which is necessary for medical point cloud registration.
@@ 1.2 Contributions
In this paper, we propose a new point-matching-based registration method for medical point cloud registration that addresses the lack of semantic information and biomechanical energy regularization in ICP. Our method first extends the semantic label-informed point-matching to non-rigid medical point cloud registration to ensure semantic-consistent matching, and then introduces a novel deformation representation in point clouds, along with explicit biomechanical energy regularization, into point cloud registration. The method was tested on three different datasets and achieved the best surface matching performance compared with other baseline algorithms. It is also tested on an additional dataset to show its ability to be integrated with upstream learning-based labelers, achieving accurate point cloud alignment. Our contributions include: (1) systematically introducing semantic label consistent point matching for non-rigid point matching-based point cloud registration; (2) a novel control-point-based deformation representation in optimization-based point cloud registration that enables practical elastic regularization; (4) validation of the effectiveness of our method in medical point clouds, with a fixed set of hyperparameters that underwent minimal tuning.
The rest of this paper is organized as follows. Section 2 describes the proposed method, and Section 3 presents validation results on four datasets, ablation studies, and sensitivity analyses. Section 4 provides the discussion.
@@ 2 Methods
We define the source point cloud as $P=\{\bm{p_{0}},\bm{p_{1}},...,\bm{p_{N}}\}$ and the target as $Q=\{\bm{q_{0}},\bm{q_{1}},...,\bm{q_{M}}\}$, with the points $\bm{p},\bm{q}\in\mathbb{R}^{3}$. The point clouds are associated with semantic labels $L_{p}=\{x_{0},x_{1},...,x_{N}\}$ and $L_{q}=\{y_{0},y_{1},...,y_{M}\}$. The labels belong to the label set $L=\{l_{1},l_{2}...,l_{K}\}$, where $K$ is the number of labels. We assume that the observed points have at least two labels, $K\geq 2$. Apart from the 3D location of the points, the point clouds also have local normals at each point: $N_{p}=\{\bm{n_{0}},...,\bm{n_{N}}\}$, $N_{q}=\{\bm{m_{0}},...,\bm{m_{M}}\}$. The local normals can be estimated using surface reconstruction methods such as ball pivoting (Bernardini et al. (2002)). The registration workflow of Semantic-ICP (SemICP), is shown in Fig. 1; it first estimates a rigid transform $\bm{T}\in SE(3)$, then refines it by non-rigid registration.
@@ 2.1 Rigid Initialization
The optimal rigid transform $\bm{T}\in SE(3)$ is iteratively estimated by minimizing a cost function that measures the difference between point clouds, as described in Algorithm 1. SemICP uses the semantic information during point pair matching, such that matched point pairs have the same semantic labels. At each iteration $i$, for each point $\bm{p_{j}}$ from $P$, we find the closest point $\bm{q_{j}}$ in $Q$ that has the smallest L2 distance with the same semantic label as $\bm{p_{j}}$. KD-tree is used to accelerate the search for matched point pairs. The update value $\Delta\bm{T}^{i}$ is automatically estimated using the autograd function in PyTorch. We use an Adam optimizer (Kingma and Ba (2014)) to minimize the cost function $L_{rigid}$ that measures the point-to-plane distance (Segal et al. (2009)) of matched point pairs at each iteration:
@@
6:    Computed closest point $\bm{q_{j}}^{i}$ with smallest L2 distance.
@@ 2.2 Non-rigid Registration
$\bm{d}_{j}$ is the surface point deformation at point $\bm{p}_{j}$. We compute $\bm{d}_{j}$ using a trilinear interpolation of the deformation at the control points that are in $\bm{p}$’s neighborhood. The neighborhood of a control point is defined as the 26 points connected with it.
@@
We use linear elastic regularization (Fischer and Modersitzki (2004); Nir et al. (2013)) to regularize the deformation. In continuum mechanics, the linear elastic energy can be described using the Navier-Lamé Equation. We approximate the Navier-Lamé Equation with the numerical differentiation at the control points and add a scaling factor to normalize the energy, where $\Delta$ is the step size of the control points.
@@
In addition to the above energy regularization, the regularization term also includes magnitude regularization $Reg_{mag}$ to reduce unwanted large deformation at the control point $\bm{c_{i}}$, and a gradient regularization $Reg_{grad}$ to encourage control grid smoothness:
@@ 3.1 Datasets
The experiments are performed on one private dataset derived from trans-oral robotic surgery (TORS), and two public datasets from Learn2Reg challenges (Hering et al. (2022)). We use the training data in the Learn2Reg challenges, since label maps for test data are not included in the public dataset link [^1] https://learn2reg.grand-challenge.org/Datasets/. We also utilize an additional private liver MR-3D US dataset that incorporates learning-based point cloud labeling and elastography for tissue stiffness measurement. 3D slicer is used to extract point clouds from the label maps (Fedorov et al. (2012)).
TORS: This dataset contains 3D freehand US and CT images of the neck collected from patients who underwent TORS for oral cancer between January 2022 and November 2024 at the Vancouver General Hospital (Vancouver, BC, Canada). This study received ethics approval from the University of British Columbia Clinical Research Ethics Board (H19-04025). Informed consent was obtained from all participants included in this study. Diagnostic CT data were collected preoperatively, and US data were collected before surgery while patients were anesthetized. A BK3500 with a 14L3 linear 2D transducer (BK Medical, Burlington, MA) and a Polaris Spectra (Northern Digital, ON, Canada) were used to collect freehand 3D US. PLUS (Lasso et al. (2014)) was used to calibrate the US probe and collect tracked US data. The image depth was set to 4 or 5 cm (depending on patient anatomy) at 9 MHz, and US imaging was carried out ipsilateral to the tumor. A medical student and an experienced research assistant labeled the carotid artery, jugular vein, and larynx/mandibular bone in 2D US frames and CT volumes, then the 2D US label maps were reconstructed in 3D. The dataset contains 24 pairs of US-CT images from 12 patients. The bifurcation point and the axis of the carotid in US and CT were used to pre-align volumes.
AbdominalMRCT (Clark et al. (2013)): The Learn2reg dataset contains 8 pairs of intra-patient abdominal magnetic resonance (MR) and CT images. The label maps include 3D segmentation of 4 anatomical labels of liver, spleen, right kidney and left kidney.
HippocampusMR (Simpson et al. (2019)): The Learn2reg challenge dataset includes MR images from 90 healthy adults and 105 adults with a non-affective psychotic disorder, along with segmentations of the head and body of the hippocampus. We use the validation registration pairs provided in the dataset’s JSON files, resulting in 60 inter-patient registration pairs.
LiverMR-3DUS (Zeng et al. (2021, 2024)): The private MR-3DUS dataset was collected from an earlier liver vibro-elastography imaging validation study with the approval of the UBC Clinical Research Ethics Board (H14-01964). The dataset contains intra-subject MR and 3D US image data from 19 healthy subjects and 20 patients with chronic liver diseases. Each subject completed MR exams with Achieva 3T or Ingenia Elition 3T X scanners (Philips Healthcare, Best, Netherlands). The abnormal image volumes were used with a voxel size of 0.9375 mm $\times$ 0.9375 mm $\times$1.5 mm. The US exam used a 3D B-mode acquisition with the EPIQ 7G system and the X6-1 xMATRIX array transducer (Philips Healthcare, Bothell, WA), with a 16 cm imaging depth and a 90 (lateral) $\times$ 60 (elevational) field of view, repeated 5 to 8 times. The voxel size for 3D US volumes was 0.468 mm $\times$ 0.3231 mm $\times$ 0.6731 mm. The dataset contains 174 MR-3D US image pairs. Gallbladder, inferior vena cava (IVC), right kidney, and major hepatic vessel branches were manually segmented by experts as registration surrogates.